OpenStudy (anonymous):

the derivative of (x)/(x+1)

OpenStudy (anonymous):

the derivative of 1/x = ln(x) right? would the derivative of x/x+1 be xln(x)?!

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Quotient_rule

OpenStudy (anonymous):

is it (x+1)-x / (x+1)^2 ?

OpenStudy (anonymous):

so the derivative of 1/x is not lnx? haha

OpenStudy (australopithecus):

you can either use product rule or quient rule assume the function is f(x) assume the neumorator as g(x) and the denominator as s(x) g(x) = x g'(x) = 1 s(x) = (x+1) s'(x) = 1 Using quient rule f'(x) = g'(x)s(x) - s'(x)g(x)/(s(x))^2 so yeah sub in the values and you have or you can use product rule assume the neumorator as g(x) and the denominator as s(x) where f(x) = x(x+1)^(-1) g(x) = x g'(x) = 1 s(x) = (x+1)^(-1) s'(x) = -(x+1) you need to use chain rule to fin the derivative of s(x) thus f'(x) = g'(x)s(x) + s'(x)g(x)

OpenStudy (anonymous):

do you know the product rule? you can look at this as x(x+1)^-1

OpenStudy (anonymous):

so is what i put incorrect?

OpenStudy (australopithecus):

No itzmashy read what I wrote

OpenStudy (anonymous):

i did haha, i'm not sure what i did wrong then

OpenStudy (anonymous):

yes, but product rule is easier to remember and the quotient rule is just a special case of the product rule

OpenStudy (australopithecus):

assume the function is f(x) assume the neumorator as g(x) and the denominator as s(x) g(x) = x g'(x) = 1 s(x) = (x+1) s'(x) = 1 f'(x) = (1(x+1) - (x)(1))/(x+1)^(2)

OpenStudy (australopithecus):

yeah I know philips :) I just have it memorized so I would use it in a case like this

OpenStudy (anonymous):

(x) = x g'(x) = 1 s(x) = (x+1) s'(x) = 1 Using quient rule f'(x) = g'(x)s(x) - s'(x)g(x)/(s(x))^2 1(x+1) - 1(x)/(x+1)^2 i simplified to (x+1) - x/ (x+1)^2 how is that different from: is it (x+1)-x / (x+1)^2 ?

OpenStudy (australopithecus):

well first of all you need to put brackets around the neumorator but they are the same otherwise

OpenStudy (anonymous):

so i did do it correctly? i thought you said it was wrong? i'm confused now

OpenStudy (australopithecus):

It is correct, there is just another method you can use on these questions involving chain rule if you know what that is?

OpenStudy (anonymous):

using chain rule u get -x(x+1)^-2, right?

OpenStudy (australopithecus):

it would give you the same answer I explained it above but I was lazy and didnt explain how I got the derivative of (x+1)^(-1)

OpenStudy (australopithecus):

oh yeah you are right :) I made a mistake

OpenStudy (anonymous):

wait.. i think i did something wrong there... lol

OpenStudy (australopithecus):

remember chain rule is g(x) = outside function d(x) = inside function chain rule is g'(d(x)) * d'(x)

OpenStudy (australopithecus):

d/dx x = 1

OpenStudy (australopithecus):

thus 1 * -(1+x)^(-2)

OpenStudy (australopithecus):

I mean you wouldn't get x because it differentiates to 1

OpenStudy (anonymous):

so am i using product rule and chain rule, or just chain rule? if its product rule then isn't it: -x(x+1)^-2+(x+1)^-1? and if its just chainrule then: -x(x+1)^-2?

OpenStudy (australopithecus):

no stop I will explain product rule lol

OpenStudy (anonymous):

i know what the product rule is, i'm just not understanding how you used it in this case

OpenStudy (australopithecus):

yeah Im not saying yu dont know what it is but i will show you how to apply it in this instance

OpenStudy (australopithecus):

so for product rule f(x) = x(x+1)^(-1) g(x) = x g'(x) = 1 s(x) = (x+1)^(-1) s'(x) = -(x+1)^(-2) we have to use chain rule to find s'(x) so lets assume d(x) is the outside function and m(x) is the inside function of s(x) so d(x) = x^(-1) d'(x) = -x^(-2) m(x) = x +1 m'(x) = 1 + 0 now chain rule is d'(m(x)) * m'(x) so we have that s'(x) = -(x+1)^(-2) * 1 = -(x+1)^(-2) thus f'(x) = g'(x)s(x) + s'(x)g(x) so the answer is 1(x+1)^(-1) + -(x+1)^(-2)(x) so I just read your question wrong sorry my brain is litterally dead from exams

OpenStudy (australopithecus):

no more confusion now

OpenStudy (anonymous):

ooh okay that makes sense now, thanks!